OpenStudy (anonymous):

Help will medal and fan directness pls not a lot of time For the transformation T, what is T-1? T : (x, y) (x + 4, y + 3) T-1: (x, y) (x - 4, y - 3) (-4x, -3y) (x + 3, y + 2)

OpenStudy (anonymous):

If P(-2, 1) is rotated 90°, its image is (2, -1) (1, -2) (-1, -2)

OpenStudy (lynfran):

using matrix 90 degree rotation is \[\left[\begin{matrix}0 & -1 \\ 1 & 0\end{matrix}\right]*\left(\begin{matrix}-2 \\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

so its c??

OpenStudy (lynfran):

we get \[\left(\begin{matrix}0*-2+(-1)*1 \\ 1*(-2)+0*1\end{matrix}\right)\]

OpenStudy (anonymous):

ok i think im confused

OpenStudy (lynfran):

\[\left(\begin{matrix}0-1 \\ -2+0\end{matrix}\right)\]

OpenStudy (anonymous):

so c? thats wat im thinking

OpenStudy (lynfran):

i just multiply out using matrix multiplication and yes \[\left(\begin{matrix}-1 \\ -2\end{matrix}\right)\]

OpenStudy (anonymous):

yay thanks.. wat about the first one? i got b but idk

OpenStudy (lynfran):

is that a minus 1 or \[t _{1}\]

OpenStudy (anonymous):

t1 i beleive

OpenStudy (lynfran):

well we.... T;(x,y) (x+4, y+3) T(1); (x,y) (x+4-1, y+3-1)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x+3) (y+2)

OpenStudy (anonymous):

?

OpenStudy (lynfran):

yep

OpenStudy (anonymous):

ohhh thank you!!!

OpenStudy (lynfran):

welcome